# Zipformer: A faster and better encoder for automatic speech recognition

### 05 Mar 2024 v3

@@ 3.5 ScaledAdam optimizer
$\bm{\Delta}_{t}=\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon},\vspace{-1.5em}$ (5)
@@
$\bm{\Delta}_{t}^{\prime}=\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon}.\vspace{-0.5em}$ (6)
@@
$\begin{split}\bm{\Delta}_{t,r}^{\prime}&=\eta\cdot\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}^{\prime}\\ &=\eta\cdot\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}.\end{split}\vspace{-0.99em}$ (7)
@@ A.1 ScaledAdam optimizer
     $\bm{\Delta}_{t}^{\prime}=\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon}$triangleright Compute parameter change.
@@
     $\bm{\Delta}_{t,r}^{\prime}\leftarrow\eta\cdot\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}$triangleright Compute parameter change by updating parameter scale.

### 10 Apr 2024 v4

@@ 3.5 ScaledAdam optimizer
$\bm{\Delta}_{t}=-\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon},\vspace{-1.5em}$ (5)
@@
$\bm{\Delta}_{t}^{\prime}=-\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon}.\vspace{-0.5em}$ (6)
@@
$\begin{split}\bm{\Delta}_{t,r}^{\prime}&=-\eta\cdot\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}^{\prime}\\ &=-\eta\cdot\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}.\end{split}\vspace{-0.99em}$ (7)
@@ A.1 ScaledAdam optimizer
     $\bm{\Delta}_{t}^{\prime}=-\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon}$triangleright Compute parameter change.
@@
     $\bm{\Delta}_{t,r}^{\prime}\leftarrow-\eta\cdot\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}$triangleright Compute parameter change by updating parameter scale.